# Lightweight Correlation-Aware Table Compression

### v1

@@ 1 Introduction
State of the Art. For this very reason, recent research proposes a change in mindset: Instead of relying only on simple encoding schemes such as RLE or DELTA, leverage correlations between columns that are present in real-world datasets [7, 8, 9, 10, 11] and represent them in a compact way; the standard encoding schemes can be applied on top. This is known as semantic or correlation-aware compression and has been shown to enable (ii). However, existing solutions for numeric columns either do not satisfy (i) or are not user-friendly: DeepSqueeze [7] is based on an auto-encoder architecture that naturally makes a column dependent on all the other columns; this incurs a slowdown during column scans and violates (i). Corra [10], while lightweight in itself, relies on user-defined functions to actually be able to compress. C3 [11] automates this process, but only applies a simple linear regression to column-pairs for its numeric columns, leaving compression potential unexplored.
Contribution. We introduce Virtual, a lightweight framework that automatically detects sparse functions inherently present in a given table, enabling both (i) fast scan times and (ii) competitive compression ratios. Virtual is designed to be a plug-in for open storage formats. In particular, we show saving factors of up to 40% over Parquet on data.gov tables and a minimal slowdown of column scans on the virtualized columns. Virtual is open source and available at https://github.com/utndatasystems/virtual.

### v3

@@ 1 Introduction
State of the Art. For this very reason, recent research proposes a change in mindset: Instead of relying only on simple encoding schemes such as RLE or DELTA, leverage correlations between columns that are present in real-world datasets [7, 8, 9, 10, 11, 12, 13] and represent them in a compact way; the standard encoding schemes can be applied on top. This is known as semantic or correlation-aware compression and has been shown to enable (ii). However, existing solutions for numeric columns either do not satisfy (i) or are not user-friendly: DeepSqueeze [7] is based on an auto-encoder architecture that naturally makes a column dependent on all the other columns; this incurs a slowdown during column scans and violates (i). Corra [12], while lightweight in itself, relies on user-defined functions to actually be able to compress. C3 [13] automates this process, but only applies a simple linear regression to column pairs for its numeric columns, leaving compression potential unexplored. Cuttle [9] and TICC [10] target more general functions defined on column-tuples, but they do not optimize for the number of reference columns that have to be loaded during scans.
Contribution. We introduce Virtual, a lightweight framework that automatically detects sparse functions inherently present in a given table, thus enabling both (i) fast scan times and (ii) competitive compression ratios. Virtual is designed to be a plug-in for open storage formats. In particular, we show saving factors of up to 40% over Parquet on data.gov tables and a minimal slowdown of column scans on the virtualized columns. Virtual is open source and available at https://github.com/utndatasystems/virtual.